OpenStudy (jackeightyone81):

(Score for Question 3: ___ of 5 points) (h) Two sidewalks in a park are represented by lines on a coordinate grid. Two points on each of the lines are shown in the tables. Sidewalk 1 x y 2 7 0 3 Sidewalk 2 x y 1 5 3 3 (a) Write the equation for Sidewalk 1 in slope-intercept form. (b) Write the equation for Sidewalk 2 in point-slope form and then in slope-intercept form. (c) Is the system of equations consistent independent, coincident, or inconsistent? Explain. (d) If the two sidewalks intersect, what are the coordinates of the point of intersection? Use the substitution m

sam (.sam.):

Are you there?

OpenStudy (jackeightyone81):

yes

OpenStudy (jackeightyone81):

answer the question its your medal

sam (.sam.):

What seems to be the problem? Where you got stucked?

OpenStudy (jackeightyone81):

i dont understand the question nor how to do it can you answer the question

sam (.sam.):

Can you rewrite the 'x' and 'y's, I'm pretty sure it's a table.

OpenStudy (jackeightyone81):

i can't that's how the question is

sam (.sam.):

I don't understand this: Sidewalk 1 x y 2 7 0 3 Sidewalk 2 x y 1 5 3 3

OpenStudy (jackeightyone81):

here i'll put a picture

sam (.sam.):

So

sam (.sam.):

The question asks you to find the equation of Sidewalk 1. To construct an equation you will need a coordinate and a slope. From the table you've got (2,7) and (0,3). Find the slope by \[m=\frac{y-y_1}{x-x_1}= \frac{7-3}{2-0}=\frac{4}{2}=2\] Now you have the slope, simply pick a coordinate and substitute the slope and the coordiate into this \[y-y_1=m(x-x_1)\] \[y-3=2(x-0) \\ \\ y=2x+3\]

sam (.sam.):

Now do question (b) using the same method

OpenStudy (jackeightyone81):

so for slope a it is y−3=2(x−0)y=2x+3 or just y=2x+3

sam (.sam.):

What do you think? of course its y=2x+3

OpenStudy (jackeightyone81):

i have no common sense at math lol

sam (.sam.):

Once you get it it's very very easy

OpenStudy (jackeightyone81):

can you answer part b im still unsure sorry for the inconvience

sam (.sam.):

No you need to try it for yourself, just follow my steps and do it

OpenStudy (jackeightyone81):

ill try